# Quasi-stationary evolution of cubic-quintic NLSE drop-like solitons in DNA-protein systems

### v1

@@ 3.1 The quasi-stationary method
The study of the perturbed cubic-quintic NLSE is crucial not only for understanding DNA-protein interactions but also for other branches of physics. In the previous analysis, the DNA-protein system was treated as an isolated system supporting the cubic-quintic NLSE. However, in reality, the DNA molecule interacts with its surrounding environment. Modeling the interaction between a single DNA molecule and its environment is a complex problem, but in a simplified scenario, it can be reduced to two primary effects: dissipation and the influence of the endogenous field [60]. For this analysis, we assume that the solvent consists of water molecules, which dampen the vibrations of the DNA bases, leading to the dissipation of nonlinear excitations in the system. To account for this, we introduce a perturbation term in Eq. (8) to study the evolution of the base-pair dynamics. To model the damping effects, we apply the multiple scales perturbation approach proposed by Y. Kodoma and M. J. Ablowitz, commonly known as the quasi-stationary method [61]. This method offers a significant advantage: it allows us to identify coherent, localized structures that emerge from the initial soliton-like solution. Before proceeding with the perturbative analysis of the molecular system, we first provide a brief overview of the quasi-stationary method.
@@
being $K$ and $F$ nonlinear functions of $q$, $q_{x},...$. We are considering that the nonlinear equation (11) depends implicitly of $\epsilon$, such as when we fix the parameter $\epsilon=0$ we obtain the unperturbed equation.
@@
where $q_{0}$ is a solitary wave or a soliton solution. We write the solution now in terms of fast and slow variables:
@@
It is denoted by $\theta_{i}(i=1,...,m)$ the fast variables and by $T=\epsilon{t}$ slow variables, and $P_{i}(i=1,...,N)$ are parameters which depend on the slow variables. Here the variables $\theta$ satisfy $\partial{\theta}/\partial{x}=1$, and $\partial{\theta}/\partial{t}=-P_{1}$ and $P_{1}=P_{1}(T)$ in order to remove secular terms. Therefore, these solutions are called quasi-stationary solutions $q=\hat{q}(\theta,T,\epsilon)$. Thus, we assume an expression of the form
@@
Substituting $\hat{q}$ in the equation (11) we come up with the first order equations:
$L(\partial_{\theta_{i}},\hat{q}_{0})\hat{q}_{1}{\equiv}F(\hat{q}_{0})-\frac{\partial{K}}{\partial{q_{t}}}\cdot{q_{T}}|_{q=\hat{q}_{0}}{\equiv}\hat{F}.$ (15)
Here $L(\partial_{\theta_{i}},\hat{q}_{0})u=0$ is a linearized equation of $K(q,q_{t},q_{x},...)=0$. We can denote by $v_{i}(i=1,...,M)$ the $M$ solutions of the homogeneous adjoint problem satisfying
@@
where $L^{A}$ is the adjoint operator to $L$, we have
@@
The last equation can be integrated to give the secularity conditions and, consequently, we will be in position to compute the solution for $\hat{q}_{1}$. by considering certain boundary conditions.
@@ 3.2 First order perturbation for the drop-soliton solution of the CQNSE
To implement the quasi-stationary method outlined above, we begin with the perturbed form of the cubic-quintic nonlinear Schrödinger equation (NLSE)
$i\varphi_{t}+\frac{1}{2}\varphi_{xx}+{\varphi}{|\varphi|^{2}}-\alpha\varphi{|\varphi|^{4}}=-i{\epsilon}\varphi.$ (18)
The right side of the equation (18) describes the damping effect caused by the viscous medium. The parameter $0<\epsilon\ll 1$ is a coupling coefficient of the strength of viscous damping, which is very small at physiological temperatures. Taking the last assumption into account and in order to study the evolution of the nonlinear excitations in the DNA molecular chain we follow the perturbative analysis described on the previous section. When $\epsilon=0$ we obtain the unperturbed equation (10), which admits the following drop soliton-like solution [62]
@@
Following the quasi-stationary method we write the soliton-like solution (19) in terms of fast and slow variables. Thus, we introduce a slow time variable $T={\epsilon}t$, and $A$, $V$, $\theta_{0}$ , $\sigma_{0}$ are functions of this time scale. We are able to write the envelope one soliton solution as
@@
Replacing (21) on (18) and assuming the quasi-stationarity we obtain
@@
We assume an expression for $\hat{\varphi}$
$\varphi=\hat{\varphi}_{0}(\theta,T)+\epsilon\hat{\varphi}_{1}+{\epsilon}^{2}\hat{\varphi}_{2}+...$
Neglecting higher order of $\epsilon$, i.e., we focus on the first order soliton perturbation, such that
$\varphi=\hat{\varphi}_{0}(\theta,T)+\epsilon\hat{\varphi}_{1},$
@@
$\hat{\varphi}_{0}=\frac{A}{\bigg{[}1+\sqrt{1-\frac{4}{3}\alpha{A^{2}}}\,\text{cosh}(\sqrt{2}A(\theta-\theta_{0})\bigg{]}^{1/2}}$
and $\hat{\varphi}_{1}=\phi_{1}+{i}\psi$, being $\phi_{1}$ and $\psi_{1}$ real functions. Substituting these expressions into the Eqs. (22) and (23), we obtain the following system
@@
The operators $L_{1}$ and $L_{2}$ are self-adjoint and $L_{1}\hat{\varphi}_{0\theta}=0$, $L_{2}\hat{\varphi}_{0}=0$. Consequently, the equations (24a) and (24b) have localized solutions around the drop soliton. The conditions of solvability of the system are given by the following secularity conditions
@@
Using equation (28) and (29) and the secularity condition (27) we obtain the following solution
@@
with $\frac{4}{3}\alpha{A^{2}}>0$ where $C_{1}$ and $C_{2}$ are arbitrary constants, with $N_{1}$, $N_{2}$, $N_{3}$ and $N_{4}$ defined as follows
@@
We construct the general solution as before, using the equations (31a) and (31b) and the variation of parameters. Thus, we have
@@
with $\frac{4}{3}\alpha{A^{2}}>0$. The expressions for $M({n})$, $n{\leq}8$ are given by
@@
Now, let’s us consider $\theta_{0}=\sigma_{0}=0$ in equation (19), which represents the center of the soliton and the center of the soliton phase, respectively, and $V_{T}=0$ from the secularity conditions (27). Thus, for (30) and (32), we have
@@
Imposing the boundary conditions $\phi_{1}|_{\theta=0}=C$ and $\phi_{{1}\theta}|_{\theta=0}=0$ we find the values for the arbitrary constants
@@
Similar for (33) we obtain that
@@
with $\phi_{1}$ and $\psi_{1}$ given by and (32). From the solution in Eq. (35), we can see that under perturbative effects, the drop soliton solution of Eq. (18) decays over time, as shown in Fig. 2. We can see from the solution (35) that under the perturbative effects the drop soliton solution of the equation (18) decay in time as shown in Fig. 2. For the case under consideration, $\alpha$ is sufficiently small, returning us to the case of the perturbed NLSE [33]. We observe that the amplitude of the soliton decreases as time progresses, a consequence of the secularity conditions, while its velocity remains constant. Additionally, from the figures, we can note that the damping effect increases over a short period when the damping coefficient is larger. In general, the perturbed term in Eq. (18) can take the form of any function $R(\varphi)$. For the specific case of DNA internal dynamics, we use $R(\varphi)=\varphi$ to model the effect of a viscous medium. A similar approach is seen in the work by A. Biswas, where a different form of $R(\varphi)$ is considered for unchirped fibers in the context of the cubic-quintic NLSE [63].
@@
4 Some comments on the transcription process of DNA transcription
@@
The above analytical results can be applied to the transcription process, which is well-known for producing mRNAs that carry the genetic information necessary for protein synthesis. In addition to mRNA, transcription also generates transfer RNA (tRNA), ribosomal RNA (rRNA), and other RNA molecules that play crucial roles in cellular structure and catalysis. All of these RNA molecules are synthesized by RNA polymerase enzymes, which transcribe a DNA sequence into an RNA copy. The enzyme binds to the promoter region of the DNA and initiates transcription at a specific start site within the promoter. It continues synthesizing the RNA strand until it reaches a termination signal, at which point both the polymerase and the newly formed RNA molecule are released. Therefore, an RNA chain of 5000 nucleotides takes about 3 minutes to complete [64]. The DNA binds to the active site of the protein, causing the base pairs to separate and unwinding the DNA double helix. Recently, numerical studies have been made to understand the RNAp-DNA dynamics taking into account the helical coupling and the inhomogeneties of DNA [38, 39].
@@
Figure 3: Schematic representation of the transcription process. The transcription process begins with the double-helix unwinding and unzipping (as in replication).
@@
In this study, we neglect these contributions due to the fact that the contribution of inhomogeneities is small compared to homogeneous stacking. Besides, in these studies the localized inhomogeneity does not affect the bubble soliton sizes, even when the amplitude of inhomogeneity increases and the number of base pairs participating in the opening does not change due to this pattern of inhomogeneities. Moreover, we are interested in the analytic evolution of soliton-like solutions in the DNA-protein system under the viscous medium. From these numerical studies [38, 39] we obtained the approximate values of the hydrogen bonds coefficient ($\mu$) and the stacking energy costant($J$), the elastic constants ($k_{1}$) of the RNAp molecule and phonon chain ($k_{2}$)) and their coupling coefficients to the DNA molecule ($\alpha_{1}$ and $\alpha_{2}$), respectively. Also, the mass of the hydrogen atom attached to the base ($m_{1}$) and the mass of the peptide ($m_{2}$) are well-known.
@@
$\alpha_{2}=0.081\ \ eV;\ \ \ \ m_{1}=1.0079\ \ u.m.a;$
$m_{2}=300\ \ u.m.a.;\ \ \ \ a=3.4\ \ $\mathrm{\SIUnitSymbolAngstrom}$$
Besides, it’s known that the cubic-quintic NLSE’s potential has several vacuum states, i.e, classical configurations with minimal energy. Therefore the solution of the equation of motion that lead to asymptotic values of $\psi(s)$ have to coincide with minimum potentials. Thus, we can find the velocities of the displacement of the peptide and the hydrogen atom are given from:
$\dfrac{A}{\rho_{0}}=-2-\dfrac{3}{4}\dfrac{k_{3}^{2}}{k_{1}k_{5}}\Bigg{(}1\pm\sqrt{1+4\dfrac{k_{1}k_{5}}{k_{3}^{2}}}\Bigg{)},$
without any loss of generality we can fix the value $\rho_{0}=1$, it can be easily demonstrated that the solution properties can be represented in combinations of $A/\rho_{0}$. Let us remind that for obtaining information on DNA properties we consider $A>1$, and in our particular case the drop condition is $\psi\to 0$ whether $x\pm\infty$. Thus, we can choose arbitrary $v_{1}$ and $v_{2}$ that satisfies this condition.
In previous studies, S. Sdravković and M. V. Satarić [60] found, using the Peyrard’s DNA model, the values of the speed and width of the soliton-like solution to be $3750\ m/s$ and $580$\mathrm{\SIUnitSymbolAngstrom}$$, respectively, which means that 170 bases in each strand is covered by the soliton. The number of base pairs that participate in the opening process depends on the nature of the soliton. However, it is advantageous to have base pair opening with participation by fewer bases.
Finally, the value of the coupling coefficient of the strength of viscous damping $\epsilon$ depends on the physiological temperature (310 K). Thus, we have that $\epsilon$ is of the order of $10^{-11}kg/s$ and the experimental value is $\epsilon\approx 6\times 10^{-11}kg/s$.

### v2

@@ 3 Perturbative analysis for the cubic-quintic NLSE solitons
As it was already mentioned in the previous discussion, the DNA-protein system was treated as an isolated system supporting the cubic-quintic NLSE. However, in reality, the DNA molecule interacts with its surrounding environment. Modelling the interaction between a single DNA molecule and its environment is a complex problem, but in a simplified scenario, it can be reduced to two primary effects: dissipation and the influence of the endogenous field [86]. For this analysis, we assume that the solvent consists of water molecules, which dampen the vibrations of the DNA bases, leading to the dissipation of nonlinear excitations in the system. To model the damping effects, we apply the multi-scales perturbation approach proposed by Y. Kodoma and M. J. Ablowitz, commonly known as the quasi-stationary method [87]. This method offers a significant advantage: it allows us to identify coherent, localized structures that emerge from the initial soliton-like solution. Before proceeding with the perturbative analysis of the molecular system, we first provide a brief overview of the quasi-stationary method.
@@ 3.1 The quasi-stationary method
being $K$ and $F$ nonlinear functions of $q$, $q_{x},...$. We consider that the nonlinear equation (12) depends implicitly of $\epsilon$, such that setting $\epsilon=0$ yields the unperturbed equation
@@
where $q_{0}$ is a solitary wave or a soliton solution. Now, we write the solution in terms of fast and slow variables:
@@
Here $\theta_{i}(i=1,...,m)$ are fast variables, $T=\epsilon{t}$ represent slow variables, and $P_{i}(i=1,...,N)$ are parameters that depend on the slow variables. Additionally, the variables $\theta$ satisfy $\partial{\theta}/\partial{x}=1$, $\partial{\theta}/\partial{t}=-P_{1}$ and $P_{1}=P_{1}(T)$ in order to remove secular terms. Therefore, these solutions are called quasi-stationary solutions $q=\hat{q}(\theta,T,\epsilon)$. Thus, we assume an expression of the form
@@
Substituting $\hat{q}$ into Eq. (12) yields the first-order equations:
$L(\partial_{\theta_{i}},\hat{q}_{0})\hat{q}_{1}{\equiv}F(\hat{q}_{0})-\frac{\partial{K}}{\partial{q_{t}}}\cdot{q_{T}}|_{q=\hat{q}_{0}}{\equiv}\hat{F},$ (16)
where $L(\partial_{\theta_{i}},\hat{q}_{0})u=0$ is a linearized equation of $K(q,q_{t},q_{x},...)=0$. We can denote by $v_{i}(i=1,...,M)$ to the $M$ solutions of the homogeneous adjoint problem satisfying
@@
with $L^{A}$ as the adjoint operator of $L$, we have
@@
If the last equation can be integrated to yield the secularity conditions, we will be in a position to compute the solution for $\hat{q}_{1}$.
@@ 3.2 First order perturbation for the drop-soliton solution of the CQNSE
To implement the quasi-stationary method outlined above to our DNA-protein system, we begin with the perturbed form of the cubic-quintic NLSE
$i\varphi_{t}+\frac{1}{2}\varphi_{xx}+{\varphi}{|\varphi|^{2}}-\alpha\varphi{|\varphi|^{4}}=-i{\epsilon}\varphi,$ (19)
which was obtained by the following scale transformation of Eq. (10)
$\psi=e^{ik_{1}t}\phi(s,t),\ \phi=\varphi(\beta s,\beta^{2}{t}),$ (20)
with $\beta=(k_{3}/2)^{1/2}$ and $\alpha=-2k_{5}/k_{3}$, we also introduce the reparametrization $s\to x$.
The right side of the equation (19) describes the damping effect caused by the viscous medium. The parameter $0<\epsilon\ll 1$ is a coupling coefficient of the strength of viscous damping, which is very small at physiological temperatures. Considering the previous assumption, and in order to investigate the evolution of nonlinear excitations in the DNA molecular chain, we employ the perturbative analysis outlined in the preceding section. When $\epsilon=0$, we recover the unperturbed cubic-quintic NLSE, which admits the following drop soliton-like solution [88]
@@
Using the quasi-stationary method, we express the soliton-like solution (21) in terms of both fast and slow variables. To this end, we introduce a slow time variable $T={\epsilon}t$, and $A$, $V$, $\theta_{0}$ , $\sigma_{0}$ being functions of this new time scale. This allows us to express the envelope of the one-soliton solution as follows:
@@
Substituting (23) into (19) and assuming quasi-stationarity, we obtain:
@@
We assume that $\hat{\varphi}$ is expressed as
$\varphi=\hat{\varphi}_{0}+\epsilon\hat{\varphi}_{1}+{\epsilon}^{2}\hat{\varphi}_{2}+...$ (26)
Neglecting higher order of $\epsilon$, we focus on the first-order perturbation, such that
$\varphi=\hat{\varphi}_{0}(\theta,T)+\epsilon\hat{\varphi}_{1},$ (27)
@@
$\hat{\varphi}_{0}(\theta,T)=\frac{A}{\bigg{[}1+\sqrt{1-\frac{4}{3}\alpha{A^{2}}}\,\text{cosh}(\sqrt{2}A(\theta-\theta_{0})\bigg{]}^{1/2}}$ (28)
and $\hat{\varphi}_{1}=\phi_{1}+{i}\psi$, being $\phi_{1}$ and $\psi_{1}$ real functions. Substituting expressions (27) and (28) into Eqs. (24) and (25), we obtain the following system of equations
@@
The operators $L_{1}$ and $L_{2}$ are self-adjoint and $L_{1}\hat{\varphi}_{0\theta}=L_{2}\hat{\varphi}_{0}=0$. Consequently, the equations (29a) and (29b) have localized solutions around the drop soliton. The conditions of solvability of the system are given by the following secularity conditions
@@
In the presence of the perturbation term the parameters deform adiabatically, these adiabatic variations are given by (32). This is an easy way to corroborate the equations obtained by the Fredholm alternative. In order to find the perturbed soliton solutions, we need to solve the homogeneous part of (29a) and (29b). Starting from (29a), we can easily see that the solutions are:
@@
Using Eqs. (33) and (34), together with the secularity condition Eq. (32), we obtain the following solution
@@
with $\frac{4}{3}\alpha{A^{2}}>0$, where $C_{1}$ and $C_{2}$ are arbitrary constants. Additionally, $N_{1}$, $N_{2}$, $N_{3}$ and $N_{4}$ are defined as follows
@@
We can derive the general solution as before by applying the method of variation of parameters, which leads to Eqs. (36a) and (36b). Consequently, the general solution for $\psi_{1}$ is found to be:
@@
with $\frac{4}{3}\alpha{A^{2}}>0$. The expressions for $M({n})$, where $n{\leq}8$, are given by
@@
Now, let us consider $\theta_{0}=\sigma_{0}=0$ in Eq. (21), which correspond to the center of the soliton and the soliton phase, respectively. Additionally, from the secularity conditions Eq. (32), we have $V_{T}=0$. Thus, for Eqs. (35) and (37), we obtain:
@@
Imposing the boundary conditions $\phi_{1}|_{\theta=0}=C$ and $\phi_{{1}\theta}|_{\theta=0}=0$, we find the values for the arbitrary constants to be
@@
Similarly, for Eq. (38), we obtain that
@@
with $\phi_{1}$ and $\psi_{1}$ given by Eq. (35) and Eq. (37), correspondingly. From the solution in Eq. (40), it is evident that under perturbative effects, the drop soliton solution of Eq. (19) decays over time, as shown in Fig. 2. We observe that the amplitude of the soliton decreases as time progresses, a consequence of the secularity conditions, while its velocity remains constant. Additionally, from the figures, we can note that the damping effect increases over a short period when the damping coefficient is larger. In general, the perturbed term in Eq. (19) can take the form of any function $R(\varphi)$. For the specific case of DNA internal dynamics, we use $R(\varphi)=\varphi$ to model the effect of a viscous medium. A similar approach is seen in the work by A. Biswas, where a different form of $R(\varphi)$ is considered for unchirped fibres in the context of the cubic-quintic NLSE [89].
@@
We can also note from our analysis that the higher-order nonlinearity lead to unusual and contrasting effects of the soliton self-spreading (or self-compression), due to their amplification (or absorption). The soliton can be compressed as it travels in an absorbing medium and it can self-spread in its adiabatic amplification. This effect, which was original studied in the context of optical fibres [79], cannot be accomplish with the standard coherent states approach once it leads to a NLSE.
To assess the impact of the viscous medium on the displacement of hydrogen bonds and peptides, we need to substitute Eq. (40) into Eq. (20), using the explicit forms given by Eqs. (6) and (7), respectively. It is expected that the quasi-stationary evolution of drop solutions, as shown in Fig. 2, will directly affect non-classical solitonic structures, such as compactons and anti-compactons. This will result in deformations of their profiles, leading to either a fast or slow decrease in amplitude, as well as compression of the soliton’s width. The latter will depend on whether the term corresponds to a loss or gain, as indicated by the sign of the viscous term. Another important, though less direct, consequence is the potential emergence of new soliton structures arising from the perturbed solution. As demonstrated in previous studies, the implementation of the quasi-stationary approach directly impacts the formation of novel structures in solitons embedded in lipid membranes [80]. However, the complex mathematical forms of the perturbed soliton in the DNA-protein system will require a more detailed analysis to confirm our main hypothesis. To ensure the robustness of our findings, this analysis will be addressed in a future work.
4 Some comments on the process of DNA transcription
@@
The analytical results discussed above can be applied to the transcription process, which is widely recognized for producing mRNA that carries the genetic information required for protein synthesis. In addition to mRNA, transcription also produces other essential RNA molecules, including transfer RNA (tRNA), ribosomal RNA (rRNA), and various non-coding RNAs, each of which plays a vital role in cellular structure and catalysis. All of these RNA molecules are synthesized by RNA polymerase enzymes, which transcribe DNA sequences into RNA copies. The enzyme binds to the promoter region of the DNA and initiates transcription at a specific start site within the promoter. Then it synthesizes the RNA strand, progressing along the DNA template until it encounters a termination signal. Upon reaching this signal, both the RNA polymerase and the newly synthesized RNA molecule are released. Therefore, the synthesis of an RNA chain consisting of 5,000 nucleotides takes approximately 3 minutes to complete [90]. The DNA binds to the active site of the protein, causing the base pairs to separate and unwinding the DNA double helix. Recently, numerical studies have been made to understand the RNAp-DNA dynamics taking into account the helical coupling and the inhomogeneties of DNA [41, 42].
@@
Figure 3: Schematic representation of the transcription process. The transcription process begins with the double-helix unwinding and unzipping (as in the replication process).
@@
In this study, we neglect contributions of inhomogeneities once that the effect of these contributions is small compared to the homogeneous stacking. Besides, in these studies the localized inhomogeneity does not affect the bubble soliton sizes, even when the amplitude of inhomogeneity increases and the number of base pairs participating in the opening does not change due to this pattern of inhomogeneities. Moreover, we are interested in the analytic evolution of soliton-like solutions in the DNA-protein system under the viscous medium. From these numerical studies [41, 42] we obtained the approximate values of the hydrogen bonds coefficient ($\mu$) and the stacking energy costant($J$), the elastic constants ($k_{1}$) of the RNAp molecule and phonon chain ($k_{2}$) and their coupling coefficients to the DNA molecule ($\alpha_{1}$ and $\alpha_{2}$), respectively. Also, the mass of the hydrogen atom attached to the base ($m_{1}$) and the mass of the peptide ($m_{2}$) are well-known.
@@
$\alpha_{2}=0.081\ \ eV;\ \ \ \ m_{1}=1.0079\ \ a.m.u;$
$m_{2}=300\ \ a.m.u.;\ \ \ \ a=3.4\ \ $\mathrm{\SIUnitSymbolAngstrom}$$
Besides, it is known that the cubic-quintic NLSE has several vacuum states, i.e, classical configurations with minimal energy. Therefore, the solution of the equation of motion that lead to asymptotic values of $\psi(s)$ in Eq. (10) has to coincide with minimum potentials [38]. Thus, we find that the velocities of the displacement of the peptide and the hydrogen atom are given by:
$\dfrac{A}{\rho_{0}}=-2-\dfrac{3}{4}\dfrac{k_{3}^{2}}{k_{1}k_{5}}\Bigg{(}1\pm\sqrt{1+4\dfrac{k_{1}k_{5}}{k_{3}^{2}}}\Bigg{)},$ (41)
without any loss of generality we fix $\rho_{0}=1$, once that the solution properties can be represented as combinations of $A/\rho_{0}$ with $A>1$. Consequently, we can choose arbitrary $v_{1}$ and $v_{2}$ that satisfies the condition given by Eq. (41).
In previous studies, S. Sdravković and M. V. Satarić [86] determined, using the Peyrard’s DNA model, that the speed and width of the soliton-like solution are $3750\ m/s$ and $580$\mathrm{\SIUnitSymbolAngstrom}$$, respectively. This implies that the soliton spans approximately 170 bases in each DNA strand. The number of base pairs involved in the opening process depends on the nature of the soliton; however, it is preferable for fewer base pairs to participate in the opening process. Finally, the value of the coupling coefficient for the strength of viscous damping, $\epsilon$, is temperature-dependent, typically around physiological temperatures ($\sim 300$ K). Therefore, $\epsilon$ is of the order of $10^{-11}kg/s$, with the experimental value being approximately $\epsilon\approx 6\times 10^{-11}kg/s$.